A. 2 3/7 is the same as.

B. If six orange are three - eighths of a bag, how many oranges are in one bag?

C. 6.35 is not equivalent to?

D. Solve for x in the proportion 3/9 = x/24

E. If A:B:C =1:3:4 and C:D = 3:4 the A:B:C:D is equal to


1
Expert's answer
2021-05-07T10:39:07-0400

A. "2\frac{3}{7} = \frac{6}{7}"

B. Proportion:

6 oranges - "\frac{3}{8}"

Total oranges -"\frac{8}{8}"

Total oranges "= \frac{6 \times 8/8}{3/8} = \frac{6 \times 8}{3} = 16"

C. 6.35 is not equivalent to "\frac{6}{35}"

D. "\frac{3}{9} = \frac{x}{24}"

"x = \frac{3 \times 24}{9} = 8"

E. A:B:C:D = 3:9:12:16